# The strong convergence phenomenon

### v1

@@ 2.5. Linearization
    For every $D\in\mathbb{N}$ and self-adjoint $P\in\mathrm{M}_{D}(\mathbb{C})\otimes\mathbb{C}\langle x_{1},\dots,x_{2r}\rangle$ of degree one,
    $\lim_{n\to\infty}\|P(\bm{U}^{N},\bm{U}^{N*})\|=\|P(\bm{u},\bm{u}^{*})\|\quad\text{in probability}.$
@@
Theorem 2.19 is due to Pisier [114, 117], but the elementary proof we present here is due to Lehner [84, §5.1]. We will need a classical lemma.
@@
For any operator $X$ in a $C^{*}$-algebra $\mathcal{A}$, define its self-adjoint dilation $\tilde{X}=e_{1}e_{2}^{*}\otimes X+e_{2}e_{1}^{*}\otimes X^{*}$ in $\mathrm{M}_{2}(\mathbb{C})\otimes\mathcal{A}$. Then $\|X\|=\|\tilde{X}\|$ and $\mathrm{sp}(\tilde{X})=-\mathrm{sp}(\tilde{X})$.
@@
We first note that $\|\tilde{X}\|^{2}=\|\tilde{X}^{2}\|=\|e_{1}e_{1}^{*}\otimes XX^{*}+e_{2}e_{2}^{*}\otimes X^{*}X\|=\|X\|^{2}$. To show that the spectrum is symmetric, it suffices to note that $\tilde{X}$ is unitarily conjugate to $-\tilde{X}$ since $U\tilde{X}U^{*}=-\tilde{X}$ with $U=(e_{1}e_{1}^{*}-e_{2}e_{2}^{*})\otimes\mathbf{1}$. ∎
@@
Now note that the $r\times r$ block matrix $\tilde{A}=(\tilde{A}_{ij}+c1_{i=j}\mathbf{1})_{i,j\in[r]}\in\mathrm{M}_{2Dr}(\mathbb{C})$ is self-adjoint, and we can choose $c$ sufficiently large so that it is positive definite. Then we may write $\tilde{A}=B^{*}B$ for $B\in\mathrm{M}_{2Dr}(\mathbb{C})$. Now view $B$ as an $1\times r$ block matrix with $2Dr\times 2D$ blocks $B_{1},\ldots,B_{r}$, so that $\tilde{A}_{ij}+c1_{i=j}\mathbf{1}=B_{i}^{*}B_{j}$. Therefore $\|X\|+rc=\|Y^{*}Y\|=\|Y\|^{2}$ with $Y=\sum_{i=1}^{r}B_{i}\otimes U_{i}$. To conclude we let $C=rc$, $D^{\prime}=2Dr$, and define $A_{i}^{\prime}$ by padding $B_{i}$ with $2D(r-1)$ zero columns. ∎
@@
    For every $D\in\mathbb{N}$ and self-adjoint $P\in\mathrm{M}_{D}(\mathbb{C})\otimes\mathbb{C}\langle x_{1},\dots,x_{r}\rangle$ of degree one,
    $\mathrm{sp}\big{(}P(\bm{X}^{N})\big{)}\subseteq\mathrm{sp}\big{(}P(\bm{x})\big{)}+o(1)[-1,1]\quad\text{with probability}\quad 1-o(1).$
@@
    For every $D\in\mathbb{N}$ and self-adjoint $P\in\mathrm{M}_{D}(\mathbb{C})\otimes\mathbb{C}\langle x_{1},\dots,x_{r}\rangle$
    $\|P(\bm{X}^{N})\|\leq\|P(\bm{x})\|+o(1)\quad\text{with probability}\quad 1-o(1).$
We omit the proof, which may be found in [60] or in [59] (see also [99, §10.3]). Let us note that while this theorem only gives an upper bound, the corresponding lower bound will often follow from Lemma 2.13.
@@
    For every $p,D\in\mathbb{N}$ and self-adjoint $P\in\mathrm{M}_{D}(\mathbb{C})\otimes\mathbb{C}\langle x_{1},\dots,x_{r}\rangle$ of degree one,
    $\lim_{N\to\infty}\mathop{\mathrm{tr}}\big{(}P(\bm{X}^{N})^{2p}\big{)}=({\mathop{\mathrm{tr}}}\otimes\tau)\big{(}P(\bm{x})^{2p}\big{)}\quad\text{in probability}.$
@@ 2.6. Positivization
The linearization trick of the previous section states that if we work with general matrix coefficients, it suffices to consider only polynomials of degree one. We now introduce (in the setting of group $C^{*}$-algebras) a complementary principle: if we admit polynomials of any degree, it suffices to consider only polynomials with positive scalar coefficients. This positivization trick due to Mikael de la Salle appears in slightly different form in [90, §6.2]. [^10] The form of this principle that is presented here was explained to the author by de la Salle.
@@
The key feature of this property is the polynomial dependence on degree $q$. This is a major improvement over the trivial bound obtained by applying the triangle inequality and Cauchy–Schwarz, which would yield such an inequality with an exponential constant |$|\{g\in\mathbf{G}:a_{g}\neq 0\}|^{1/2}\leq(2r+1)^{q/2}$.
@@
    For every self-adjoint $P\in\mathbb{R}_{+}\langle x_{1},\dots,x_{2r}\rangle$
    $\|P(\bm{U}^{N},\bm{U}^{N*})\|\leq\|P(\bm{u},\bm{u}^{*})\|+o(1)\quad\text{with probability}\quad 1-o(1).$
@@
Alternatively, when the positivization trick is applied to the polynomial method, it is possible to apply a variant of the argument directly to the limiting object that appears in the proof, avoiding the need to invoke properties of the random matrices. This form of the positivization trick is developed in [90, §6.2] (cf. Remark 3.11).
@@ 3.1.2. Asymptotic expansion
Recall that a linear functional $\nu$ on $C^{\infty}(\mathbb{R})$ is called a compactly supported (Schwartz) distribution (see [72, Chapter II]) if
@@ 3.1.3. The infinitesimal distribution
Fix $\varepsilon>0$ and a self-adjoint $P$ with positive coefficients. Moreover, let $h$ be a nonnegative smooth function that vanishes in a neighborhood of $[-\|P(\bm{u},\bm{u}^{*})\|,\|P(\bm{u},\bm{u}^{*})\|]$ and such that $h(x)=1$ for $|x|\geq\|P(\bm{u},\bm{u}^{*})\|+\varepsilon$.
@@ 3.2. Polynomial encoding
11$\scriptstyle$ 1122$\scriptstyle$ 2233$\scriptstyle$ 3311$\scriptstyle$ $1$$\scriptstyle$ 1122$\scriptstyle$ $2$$\scriptstyle$ 22
@@
Define a polynomial of degree $r(q-1)$ by
@@
We can now read off the first terms in the $\frac{1}{N}$-expansion. Recall that $g\in\mathbf{F}_{r}\backslash\{e\}$ is called a proper power if $g=v^{k}$ for some $v\in\mathbf{F}_{r}$, $k\geq 2$, and is called a non-power otherwise. Every $g\neq e$ can be written uniquely as $g=v^{k}$ for a non-power $v$.
@@ 3.3. The master inequality
by the triangle inequality, as $\nu_{0}=\mu_{P(\bm{u},\bm{u}^{*})}$ is supported in $[-K,K]$, and using the bound on $\nu_{1}(h)$. The conclusion follows using $1<\frac{Cq^{2}}{N}$. ∎
@@ 3.5. The infinitesimal distribution
It is also possible, however, to apply a variant of the positivization trick directly to $\nu_{1}$. This argument [90, §6.2] shows that the validity of Lemma 3.10 for polynomials with positive coefficients already implies its validity for all self-adjoint polynomials (even with matrix coefficients), so that the polynomial method can be applied directly to general polynomials. The advantage of this approach is that it yields much stronger quantitative bounds than can be achieved by applying Lemma 2.26. Since we have not emphasized the quantitative features of the polynomial method in our presentation, we do not develop this approach further here.
@@ 3.6.2. A new phenomenon
There is of course no contradiction between these observations: if we choose $h(x)=x^{p}$ in (3.4), then $\|h\|_{L^{\infty}[-K,K]}=K^{p}\geq e^{Cp}\|X_{\rm F}\|^{p}$ and we recover the exponential dependence on degree that was observed in Example 3.12. On the other hand, (3.4) shows that the dependence on the degree becomes polynomial when $h$ is uniformly bounded on the interval $[-K,K]$. Thus the polynomial method reveals an unexpected cancellation phenomenon that happens when the moments are combined to form bounded test functions $h$.
@@ 4.1. The free model
To work with Gaussian random matrices, we must recall how to compute moments of independent standard Gaussians $\bm{g}=(g_{1},\ldots,g_{r})$: given any $k_{1},\ldots,k_{n}\in[r]$, the Wick formula [106, Theorem 22.3] states that
@@
by the Wick formula. Consider first a noncrossing pairing $\pi$. Since pairs cannot cross, there must be an adjacent pair $\{i,i+1\}\in\pi$, and if this pair is removed we obtain a noncrossing pairing of $[n]\backslash\{i,i+1\}$. As $\mathbf{E}[G^{N}_{k_{i}}G^{N}_{k_{j}}]=1_{k_{i}=k_{j}}\mathbf{1}$, [^13] This follows from a simple explicit computation using the following characterization of GUE matrices: $G_{i}^{N}$ is a self-adjoint matrix whose entries above the diagonal are i.i.d. complex Gaussians and entries on the diagonal are i.i.d. real Gaussians with mean zero and variance $\frac{1}{N}$. we obtain
@@
On the other hand, if $\bm{\tilde{G}}^{N}$ is an independent copy of $\bm{G}^{N}$, we can computefootnote 13
[…]
@@
Free semicircular families can be constructed in various ways, guaranteeing their existence; see, e.g. [106, pp. 102–108]. Lemma 4.1 states that a family $\bm{G}^{N}$ of independent GUE matrices converges weakly to a free semicircular family $\bm{s}$.
@@
[…]
@@
[…]
@@
The final claim follows as $-\frac{1}{\pi}\mathrm{Im}\,(x+i\varepsilon)^{-1}=\frac{1}{\pi}\frac{\varepsilon}{x^{2}+\varepsilon^{2}}=\rho_{\varepsilon}(x)$ is the density of the Cauchy distribution with scale $\varepsilon$, so that $-\frac{1}{\pi}\mathrm{Im}[\mathop{\mathrm{tr}}G(x+i\varepsilon)]$ is the density of the convolution $\mu_{X_{\rm free}}*\rho_{\varepsilon}$ which converges weakly to $\mu_{X_{\rm free}}$ as $\varepsilon\to 0$. ∎
@@
These bounds admit a simple direct proof [115, p. 208].
@@ 4.2.1. The interpolation method
$\frac{d}{dq}\mathbf{E}[\mathop{\mathrm{tr}}{(X_{q}^{N})^{2p}}]=\text{sum of terms of the form}\\ \mathbf{E}\big{[}{\mathop{\mathrm{tr}}H_{a}^{N}\,(X_{q}^{N})^{m_{1}}\,\tilde{H}_{b}^{N}\,(X_{q}^{N})^{m_{2}}\,H_{a}^{N}\,(\tilde{X}_{q}^{N})^{m_{3}}\,\tilde{H}_{b}^{N}\,(\tilde{X}_{q}^{N})^{m_{4}}}\big{]}\\ \phantom{\sum}\text{with }m_{1}+m_{2}+m_{3}+m_{4}=2p-4\text{ and }a,b\in\{0,1\},$
@@ 4.2.2. The crossing inequality
$w(H,\tilde{H})^{4}=\sup_{\|M_{1}\|,\|M_{2}\|,\|M_{3}\|\leq 1}\big{\|}\mathbf{E}\big{[}H\,M_{1}\,\tilde{H}\,M_{2}\,H\,M_{3}\,\tilde{H}\big{]}\big{\|}$
@@
for any random matrix $H$. If it were to be the case that $w(H,\tilde{H})$ is monotone as a function of $\mathrm{Cov}(H)$ and $\mathrm{Cov}(\tilde{H})$, then one could bound
$w(H,\tilde{H})^{4}\mathop{\stackrel{{\scriptstyle?}}{{\leq}}}N^{2}\,\|\mathrm{Cov}(H)\|\,\|\mathrm{Cov}(\tilde{H})\|\,w(G,\tilde{G})=\|\mathrm{Cov}(H)\|\,\|\mathrm{Cov}(\tilde{H})\|$
@@
Unfortunately, $w(H,\tilde{H})$ is not monotone as a function of $\mathrm{Cov}(H)$ and $\mathrm{Cov}(\tilde{H})$, so the above reasoning does not apply directly. However, we can use a trick to rescue the argument. The key observation is that the parameter $w(H,\tilde{H})$ can be “symmetrized” by applying the Cauchy–Schwarz inequality as is illustrated informally in Figure 4.2. This results in two symmetric terms—a single pair which is readily bounded by $\|\mathbf{E}[H^{2}]\|$, and a double crossing that is a positive functional of (and hence monotone in) $\mathrm{Cov}(H)$. We can thus apply the above logic to the double crossing to replace $H$ by a GUE matrix, which yields a factor $\|\mathrm{Cov}(H)\|$. The term that remains can now be bounded using a similar argument.
@@
The matrix alignment parameter $w(H,\tilde{H})$ that appears in the proof of Lemma 4.7 (as well as the use of the Riesz–Thorin theorem in this context) was first introduced in the work of Tropp [125], which predates the discovery of the intrinsic freeness principle. Let us briefly explain how it appears there.
The idea of [125] is to mimic the classical proof of the Schwinger-Dyson equation for GUE matrices, see, e.g., [55, Chapter 2], in the context of a general Gaussian random matrix. Tropp observed that the error term that arises from this argument can be naturally bounded by $w(H,\tilde{H})$, and that this parameter is small in some examples (e.g., for matrices with independent entries).

### v2

@@ 2.5. Linearization
    For every $D\in\mathbb{N}$ and self-adjoint $P\in\mathrm{M}_{D}(\mathbb{C})\otimes\mathbb{C}\langle x_{1},\dots,x_{2r}\rangle$ of degree one, lim_n→∞ P(U^N,U^N*) = P(u,u^*)in probability.
@@
Theorem 2.19 is due to Pisier [116, 119], but the elementary proof we present here is due to Lehner [85, §5.1]. We will need a classical lemma.
@@
For any operator $X$ in a $C^{*}$-algebra $\mathcal{A}$, define its self-adjoint dilation $\tilde{X}=e_{1}e_{2}^{*}\otimes X+e_{2}e_{1}^{*}\otimes X^{*}$ in $\mathrm{M}_{2}(\mathbb{C})\otimes\mathcal{A}$. Then $\|X\|=\|\tilde{X}\|$ and $sp(X~)\mathrm${sp}($\tilde${X})=-$\mathrm${sp}($\tilde${X}).
@@
We first note that $\|\tilde{X}\|^{2}=\|\tilde{X}^{2}\|=\|e_{1}e_{1}^{*}\otimes XX^{*}+e_{2}e_{2}^{*}\otimes X^{*}X\|=\|X\|^{2}$. To show that the spectrum is symmetric, it suffices to note that $\tilde{X}$ is unitarily conjugate to $\tilde{X}$ since $U\tilde{X}U^{*}=-\tilde{X}$ with $U=(e_{1}e_{1}^{*}-e_{2}e_{2}^{*})\otimes\mathbf{1}$. ∎
@@
Now note that the $r\times r$ block matrix $\tilde{A}=(\tilde{A}_{ij}+c1_{i=j}\mathbf{1})_{i,j\in[r]}\in\mathrm{M}_{2Dr}(\mathbb{C})$ is self-adjoint, and we can choose $c$ sufficiently large so that it is positive definite. Then we may write $\tilde{A}=B^{*}B$ for $B\in\mathrm{M}_{2Dr}(\mathbb{C})$. Now view $B$ as an $1\times r$ block matrix with $2Dr\times 2D$ blocks $B_{1},\ldots,B_{r}$, so that $\tilde{A}_{ij}+c1_{i=j}\mathbf{1}=B_{i}^{*}B_{j}$. Therefore $\|X\|+rc=\|Y^{*}Y\|=\|Y\|^{2}$ with $Y=\sum_{i=1}^{r}B_{i}\otimes U_{i}$. To conclude we let $C=rc$, $D^{\prime}=2Dr$, and define $A_{i}^{\prime}$ by padding $B_{i}$ with 2D(r1)2D(r-1) zero columns. ∎
@@
    For every $D\in\mathbb{N}$ and self-adjoint $P\in\mathrm{M}_{D}(\mathbb{C})\otimes\mathbb{C}\langle x_{1},\dots,x_{r}\rangle$ of degree one, sp(P(X^N)) ⊆sp(P(x)) + o(1) [-1,1]  with probability 1-o(1).
@@
    For every $D\in\mathbb{N}$ and self-adjoint $P\in\mathrm{M}_{D}(\mathbb{C})\otimes\mathbb{C}\langle x_{1},\dots,x_{r}\rangle$ P(X^N) ≤∥P(x)+o(1)  with probability 1-o(1).
We omit the proof, which may be found in [60] or in [59] (see also [101, §10.3]). Let us note that while this theorem only gives an upper bound, the corresponding lower bound will often follow from Lemma 2.13.
@@
    For every $p,D\in\mathbb{N}$ and self-adjoint $P\in\mathrm{M}_{D}(\mathbb{C})\otimes\mathbb{C}\langle x_{1},\dots,x_{r}\rangle$ of degree one, lim_N→∞ tr(P(X^N)^2p) = (trτ)(P(x)^2p)in probability.
@@ 2.6. Positivization
The linearization trick of the previous section states that if we work with general matrix coefficients, it suffices to consider only polynomials of degree one. We now introduce (in the setting of group $C^{*}$-algebras) a complementary principle: if we admit polynomials of any degree, it suffices to consider only polynomials with positive scalar coefficients. This positivization trick was introduced in the work of Magee and de la Salle [91, §6.2]. [^10] This idea appears in [91] in a slightly different context, cf. Remark 2.27. The form of the positivization trick that is presented here was explained to the author by Mikael de la Salle.
@@
The key feature of this property is the polynomial dependence on degree $q$. This is a major improvement over the trivial bound obtained by applying the triangle inequality and Cauchy–Schwarz, which would yield such an inequality with an exponential constant $|\{g\in\mathbf{G}:a_{g}\neq 0\}|^{1/2}\leq(2r+1)^{q/2}$.
@@
    For every self-adjoint $P\in\mathbb{R}_{+}\langle x_{1},\dots,x_{2r}\rangle$ P(U^N,U^N*) ≤∥P(u,u^*) + o(1)  with probability 1-o(1).
@@
Alternatively, when the positivization trick is applied to the polynomial method, it is possible to apply a variant of the argument directly to the limiting object that appears in the proof, avoiding the need to invoke properties of the random matrices. This form of the positivization trick is developed in [91, §6.2] (cf. Remark 3.11).
@@ 3.1.2. Asymptotic expansion
Recall that a linear functional $\nu$ on $C^{\infty}(\mathbb{R})$ is called a compactly supported (Schwartz) distribution (see [73, Chapter II]) if
@@ 3.1.3. The infinitesimal distribution
Fix $\varepsilon>0$ and a self-adjoint $P$ with positive coefficients. Moreover, let $h$ be a nonnegative smooth function that vanishes in a neighborhood of $\|P(\bm{u},\bm{u}^{*})\|,\|P(\bm{u},\bm{u}^{*})\|]$ and such that $h(x)=1$ for $|x|\geq\|P(\bm{u},\bm{u}^{*})\|+\varepsilon$.
@@ 3.2. Polynomial encoding
$\scriptstyle 1$$\scriptstyle 2$$\scriptstyle 3$$\scriptstyle 1$$\scriptstyle 1$$\scriptstyle 2$$\scriptstyle 2$
@@
Define a polynomial of degree r(q1)r(q-1) by
@@
[…]
@@ 3.3. The master inequality
by the triangle inequality, as $\nu_{0}=\mu_{P(\bm{u},\bm{u}^{*})}$ is supported in [K,K][-K,K], and using the bound on $\nu_{1}(h)$. The conclusion follows using $1<\frac{Cq^{2}}{N}$. ∎
@@ 3.5. The infinitesimal distribution
It is also possible, however, to apply a variant of the positivization trick directly to $\nu_{1}$. This argument [91, §6.2] shows that the validity of Lemma 3.10 for polynomials with positive coefficients already implies its validity for all self-adjoint polynomials (even with matrix coefficients), so that the polynomial method can be applied directly to general polynomials. The advantage of this approach is that it yields much stronger quantitative bounds than can be achieved by applying Lemma 2.26. Since we have not emphasized the quantitative features of the polynomial method in our presentation, we do not develop this approach further here.
@@ 3.6.2. A new phenomenon
There is of course no contradiction between these observations: if we choose $h(x)=x^{p}$ in (3.4), then $\|h\|_{L^{\infty}[-K,K]}=K^{p}\geq e^{Cp}\|X_{\rm F}\|^{p}$ and we recover the exponential dependence on degree that was observed in Example 3.12. On the other hand, (3.4) shows that the dependence on the degree becomes polynomial when $h$ is uniformly bounded on the interval [K,K][-K,K]. Thus the polynomial method reveals an unexpected cancellation phenomenon that happens when the moments are combined to form bounded test functions $h$.
@@ 4.1. The free model
To work with Gaussian random matrices, we must recall how to compute moments of independent standard Gaussians $\bm{g}=(g_{1},\ldots,g_{r})$: given any $k_{1},\ldots,k_{n}\in[r]$, the Wick formula [108, Theorem 22.3] states that
@@
[…]
@@
On the other hand, if $\bm{\tilde{G}}^{N}$ is an independent copy of $\bm{G}^{N}$, we can computefootnote˜13
$\mathbf{E}\big[G_{k_{i}}^{N}\,A\,\tilde{G}_{k_{l}}^{N}\,B\,G_{k_{j}}^{N}\,C\,\tilde{G}_{k_{m}}^{N}\big]=\frac{1}{N^{2}}\,CBA\,1_{k_{i}=k_{j}}\,1_{k_{l}=k_{m}}$ (4.1)
@@
Free semicircular families can be constructed in various ways, guaranteeing their existence; see, e.g. [108, pp. 102–108]. Lemma 4.1 states that a family $\bm{G}^{N}$ of independent GUE matrices converges weakly to a free semicircular family $\bm{s}$.
@@
The variables $s_{i}$ are called “semicircular” because their moments $\tau(s_{i}^{p})=|\mathrm{NC}_{2}[p]|=\int_{-2}^{2}x^{p}\cdot\frac{1}{2\pi}\sqrt{4-x^{2}}\,dx$ are the moments of the semicircle distribution. Thus Lemma 4.1 recovers the classical fact that the empirical spectral distribution of a GUE matrix converges to the semicircle distribution.
@@
and $\int f\,d\mu_{X_{\rm free}}=-\frac{1}{\pi}\lim_{\varepsilon\downarrow 0}\int f(x)\,\mathrm{Im}[\mathop{\mathrm{tr}}G(x+i\varepsilon)]\,dx$ for all $f\in C_{b}(\mathbb{R})$.
@@
The final claim follows as $\frac{1}{\pi}\mathrm{Im}\,(x+i\varepsilon)^{-1}=\frac{1}{\pi}\frac{\varepsilon}{x^{2}+\varepsilon^{2}}=\rho_{\varepsilon}(x)$ is the density of the Cauchy distribution with scale $\varepsilon$, so that $\frac{1}{\pi}\mathrm{Im}[\mathop{\mathrm{tr}}G(x+i\varepsilon)]$ is the density of the convolution $\mu_{X_{\rm free}}*\rho_{\varepsilon}$ which converges weakly to $\mu_{X_{\rm free}}$ as $\varepsilon\to 0$. ∎
@@
These bounds admit a simple direct proof [117, p. 208].
@@ 4.2.1. The interpolation method
[…]
@@ 4.2.2. The crossing inequality
$w(H,\tilde{H})^{4}=\sup_{\|M_{1}\|,\|M_{2}\|,\|M_{3}\|\leq 1}\big\|\mathbf{E}\big[H\,M_{1}\,\tilde{H}\,M_{2}\,H\,M_{3}\,\tilde{H}\big$$\big\|$
@@
[…]
$Cov(H~)w(H,\tilde{H})^{4}\mathop{\stackrel{{\scriptstyle?}}{{\leq}}}N^{2}\,\|\mathrm{Cov}(H)\|\,\|\mathrm{Cov}(\tilde{H})\|\,w(G,\tilde{G})$$\|\mathrm$$\|\,\|\mathrm{Cov}(\tilde{H})\|$
@@
$\scriptstyle M_{1}$$\scriptstyle M_{2}$$\scriptstyle M_{3}$$\scriptstyle H$$\scriptstyle\tilde{H}\leq$$\frac{1}{2}$$\scriptstyle M_{1}$$\scriptstyle M_{1}^{*}$$\scriptstyle H$$\frac{1}{2}$$\scriptstyle M_{3}^{*}$$\scriptstyle M_{2}^{*}$$\scriptstyle M_{2}$$\scriptstyle M_{3}$$\scriptstyle\tilde{H}$$\scriptstyle\tilde{H}$$\scriptstyle H$
@@
The matrix alignment parameter w(H,H~)w($\tilde{H}$) that appears in the proof of Lemma 4.7 (as well as the use of the Riesz–Thorin theorem in this context) was first introduced in the work of Tropp [127], which predates the discovery of the intrinsic freeness principle. Let us briefly explain how it appears there.
The idea of [127] is to mimic the classical proof of the Schwinger-Dyson equation for GUE matrices, see, e.g., [55, Chapter 2], in the context of a general Gaussian random matrix. Tropp observed that the error term that arises from this argument can be naturally bounded by w(H,H~)w($\tilde{H}$), and that this parameter is small in some examples (e.g., for matrices with independent entries).